Q13. Super sounds has decided to reduce the annual dividends by 20% a year for the next three years. After that they will maiHello, how do I find the answer to this question?

Thank you!
Answer #1

Answer:

Correct answer is:

b. $7.36

Explanation:

D0 = $2.25

D1 = 2.25 * (1 - 20%) = $1.80

D2 = 1.80 * (1 - 20%) = $1.44

D3 = 1.44 * (1 - 20%) = $1.152

D4 = $1

Hence:

P3 = Constant dividend / Required return = 1 / 16% = $6.25

P0 = 1.80/(1 + 16%) + 1.44 ./ (1 + 16%)^2 + (1.152 + 6.25) / (1 + 16%)^3

= $7.36

As such option b is correct and other options a, c, and d are incorrect.